# CHaRNet: Conditioned Heatmap Regression for Robust Dental Landmark Localization

### v4

@@ 6 Discussion
The results demonstrate the effectiveness of the proposed CHaR module in improving 3D landmark detection for point cloud-based models, particularly in challenging scenarios involving incomplete dental models. Across multiple architectures, CHaR-augmented models consistently outperform their baseline counterparts on both MEDE and MSR, underscoring the impact of enhancing the encoding of local geometric information. This improvement is especially evident in cases with missing teeth, where baseline models often fail to capture the subtle cues necessary for accurate landmark location.
@@
When comparing performance across different dentition types, all models perform better on typical dental models than on incomplete ones. For instance, CHarNet achieves a MEDE of 0.51mm and an MSR of 87.06% on typical models, highlighting its ability to handle simpler cases with precision. However, macroaveraged results that include more complex scenarios exhibit a higher error (3.94mm MEDE) and lower success ratio (68.95% MSR), suggesting that the absence of certain teeth disrupts key structural patterns used for landmark detection. Although the CHaR module significantly narrows this performance gap, there remains a clear need for methods better equipped to handle incomplete geometries.
Among the point cloud learning models evaluated, PointMLP demonstrates the most pronounced gains with the addition of CHaR module, indicating that PointMLP’s architecture may be particularly compatible with CHaR’s approach to capturing and processing local features. Other architectures such as PointNet++ and PointNeXt also benefit from the inclusion of the CHaR module, but do not reach the same level of improvement. This disparity may be attributed to differences in how each model balances local and global feature extraction.
Qualitative analyses reinforce these quantitative findings. Visual inspections, as shown in Figure 8, reveal that while baseline models frequently misplace landmarks in regions with missing teeth, CHaR-augmented models are more robust, even when direct geometric cues are limited. The inferred landmark placements in these complex regions highlight the CHaR module’s capability to reason about plausible geometries in the face of missing data.
@@
Despite these promising results, certain limitations suggest directions for future work. Although the CHaR module significantly enhances performance in incomplete cases, current MEDE and MSR values leave room for improvement. Approaches that combine the CHaR module with additional processing steps or other network modules could be explored to address more severe or irregular forms of dental model incompleteness. Furthermore, because the CHaR module appears to rely on specific architectural features of PointMLP, efforts to generalize it to a broader range of network designs could expand its utility. Finally, while the focus here is on landmark detection accuracy, real-time processing and scalability to larger point clouds remain open questions that merit further investigation.
@@ 7 Conclusion
This work introduces the CHaRNet and the CHaR module, a novel approach to enhance 3D landmark detection in point cloud-based models, particularly for dental applications. The results demonstrate that the CHaR module significantly improves performance across multiple architectures, as evidenced by consistent reductions in MEDE and increases in MSR for both typical dental models and cases involving missing teeth. By effectively grouping landmarks associated with each tooth, the CHaR module enables the model to adapt dynamically based on the tooth’s presence or absence, assigning landmarks to a null point if the tooth is missing or accurately localizing them within the dentition if the tooth is present. This capability addresses a key challenge in dental landmark detection, where structural irregularities and missing teeth often hinder model performance.
The superior performance of the CHaR-augmented version in every model, particularly in PointMLP, highlights its suitability for incorporating the proposed module, achieving state-of-the-art results with a MEDE of 0.51 mm in typical dental models and a robust performance with 1.28 mm across all dentition types. These findings validate the efficacy of the CHaR module in improving the accuracy and robustness of landmark detection, particularly in challenging scenarios.
To contribute to the advancement of research in this domain, we are making the dataset and the code used in this study publicly available. This effort aims to address the lack of "open spirit" that is prevalent in fields such as orthodontics, in contrast to the artificial intelligence community, where openness and collaboration are common. We hope that this initiative will encourage others in the orthodontic community to share their datasets, allowing fair comparisons between AI methods, attracting more practitioners to the field of computer-assisted orthodontics, and fostering the development of new domain-specific technologies.

### v5

@@ 6 Discussion
The results demonstrate the effectiveness of the proposed CHaRM and CHaR module in improving 3D landmark detection for point cloud-based models, particularly in challenging scenarios involving incomplete dental models. Across multiple architectures, CHaR-based models consistently outperform their baseline counterparts on both MEDE and MSR, underscoring the impact of enhancing the encoding of local geometric information. This improvement is especially evident in cases with missing teeth, where baseline models often fail to capture the subtle cues necessary for accurate landmark localization. When comparing performance across different dentition types, all models perform better on standard dental models than on incomplete ones. For instance, CHaRNet (CHaR-based PointMLP) achieves a MEDE of 0.56 mm and an MSR of 85.2% on standard models, highlighting its ability to handle simpler cases with precision. However, macro-averaged results that include more complex scenarios exhibit a higher error (1.12 mm MEDE) and lower success ratio (68.5% MSR), suggesting that the absence of certain teeth disrupts key structural patterns used for landmark detection. Although the CHaR module significantly narrows this performance gap, there remains a clear need for methods better equipped to handle incomplete geometries. Qualitative analyses reinforce the superiority of CHaR-based models. Visual inspections, as shown in Figure 8, reveal that while baseline models frequently misplace landmarks in regions with missing teeth, CHaR-based models are more robust, even when direct geometric cues are limited. The inferred landmark placements in these complex regions highlight the CHaR module’s capability to reason about plausible geometries in the face of missing data.
When compared with SOTA methods, TSMDL and ALIIOS, CHaRNet delivers significant improvements in three key aspects. (i) Better teeth classification is achieved, which is critical for visually acceptable results in complex dentition where segmentation errors could catastrophically hurt individual tooth landmarks placement, as shown in Figure 9(b). (ii) Better landmark localization is achieved, mainly because of the superiority of PointMLP as point cloud feature extractor and the refinement of heatmaps’ outputs with CHaR module. (iii) Far better inference time efficiency, reducing $711\times$ and $4.5\times$ times ALIIOS and TSMDL latency, respectively, in CPU, and $102\times$ and $14.8\times$ times ALIIOS and TSMDL latency, respectively, in GPU. In absolute terms, 0.86 and 0.24 seconds in CPU and GPU, respectively, allow CHaRNet to be used in real-time clinical dental applications or processes.
@@
Despite these promising results, certain limitations suggest directions for future work. Although the CHaR module significantly enhances performance in complex cases, current MEDE and MSR values leave room for improvement. Approaches that combine the CHaR module with additional processing steps or other network modules could be explored to address more severe or irregular forms of dental model incompleteness. Furthermore, because the CHaR module appears to rely on specific architectural features of PointMLP, efforts to generalize it to a broader range of network designs could expand its utility. Finally, while the focus here is on landmark detection accuracy, real-time processing and scalability to larger point clouds and dataset sizes remain open questions that merit further investigation.
@@ 7 Conclusion
This work introduces CHaRM, a novel approach to enhance 3D landmark detection in point cloud-based models, particularly for dental applications. The results demonstrate that the CHaR module significantly improves performance across multiple point cloud architectures, as evidenced by consistent reductions in MEDE and increases in MSR for both standard dental models and cases involving missing teeth. By effectively grouping landmarks associated with each tooth, the CHaR module enables the model to adapt dynamically based on the tooth’s presence or absence, assigning landmarks to a null point if the tooth is missing or accurately localizing them within the dentition if the tooth is present. This capability addresses a key challenge in dental landmark detection, where structural irregularities and missing teeth often hinder model performance.
The superior performance of the CHaR-based version in every model, particularly in PointMLP, highlights its suitability for incorporating the proposed module, achieving SOTA results with a MEDE of 0.56 mm in standard dental models and a robust performance of 1.12 mm across all dentition types. Visual findings also validate the efficacy of the CHaR module in improving the accuracy and robustness of landmark detection, particularly in challenging scenarios. These SOTA results are achieved by eliminating the segmentation step typically found in the literature, which allowed a drastic inference time reduction, making CHaRNet suitable for real-time applications.
To contribute to the advancement of research in this domain, we are making the dataset (IOSLandmarks-1k) and the code used in this study publicly available. This effort aims to address the lack of "open spirit" that is prevalent in fields such as orthodontics, in contrast to the artificial intelligence community, where openness and collaboration are common. We hope that this initiative will encourage others in the orthodontic community to share their datasets, allowing fair comparisons between AI methods, attracting more practitioners to the field of computer-assisted orthodontics, and fostering the development of new domain-specific technologies.